# Learning from Mistakes: Enhancing Harmful Meme Detection via Misjudgment Risk Patterns

### v1

@@ 3.2.1. Bipartite Candidate Retrieval
Specifically, we use the target sample’s image vector $v_{\text{test}}^{\text{img}}$ and text vector $v_{\text{test}}^{\text{txt}}$ to separately query the subsets of samples corresponding to benign and harmful ground-truth labels. Within each labeled subset, we retrieve the top-$N$ most similar candidates from the image modality and the top-$N$ from the text modality. Subsequently, we take the union of all these retrieved candidates to form an initial candidate pool. This pool, containing samples that offer a balanced perspective on potential risks, is then passed to the reranking part for a more refined similarity assessment and re-ranking.
@@ 3.2.2. Hierarchical Harm Tree-based Reranking
The product term $\prod_{i=1}^{k}M_{i}$ ensures that the bonus $\gamma$ is awarded only if all levels from $L_{1}$ to $L_{k}$ have successfully matched, thus enforcing the top-down sequential mechanism. Consequently, the tree similarity score for a candidate, denoted as $S_{\mathcal{T}}$, is hierarchical bonuses accumulated from all continuously matched levels (top-down).
@@
where $\alpha$ is a hyperparameter that balances the importance of the image and text modalities, and $S_{\mathcal{T}}$ represents the score derived from the structured harm matching. All candidates are sorted in descending order based on their $S_{\text{final}}$ score, from which we select the top-$K$ samples, ensuring a balanced selection by choosing half from the benign candidates and the other half from the harmful ones. The corresponding misjudgment risk patterns of these selected samples are subsequently used to guide the follow-up detection.
@@ 3.3. Pattern-Augmented Reasoning for Detection
We present pattern-augmented prompt construction, where the prompt is built upon a structured template designed to orchestrate the MLLM’s analytical workflow. The template first grounds the MLLM by clearly stating the classification task and providing an objective definition of the target harm category adopted from (Lin et al., 2025). Its core component is the Critical Reflection References section, which contains the retrieved misjudgment patterns categorized based on their ground-truth labels. Patterns associated with harmful memes are aggregated and framed as Miss Reasons to alert the MLLM to subtle harmful patterns it might otherwise overlook. Conversely, patterns from benign memes are presented as False Alarm Reasons to prevent the MLLM from misclassifying non-hateful content. To ensure this guidance is used for critical reflection rather than simple imitation, the template also enforces a structured reasoning procedure that compels the model to form an independent analysis before consulting the provided support. Through this strategy, a unique and highly relevant prompt is dynamically generated for each target sample (See Appendix C Figure 8 for the detailed prompt templates).
@@ 4. Experiment Design
  • RQ2: What is the contribution of components to the harmful meme detection? We aim to analyze the effectiveness of our design, by comparing the performances of PatMD and variants of removing various components.
@@ 4.1. Datasets
PromptHarm 64.05 51.57 69.30 67.81 61.88 57.89 58.51 55.08 61.71 53.44 63.28 57.16
Pro-Cap 52.40 51.80 61.56 58.57 48.56 47.15 54.90 54.62 57.06 56.83 54.90 53.79
CapAlign 63.70 51.60 68.60 66.81 61.51 54.76 58.08 52.67 60.77 51.04 62.53 55.38
MemeAgent 68.40 62.73 73.10 73.04 62.18 58.96 70.11 69.97 63.23 61.25 67.40 65.19
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.94$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+17.95$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.53$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.20$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.10$}
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.40$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-2.61$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.00$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.17$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.49$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.71$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.48$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.72$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$}
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.82$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+15.78$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.18$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+21.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+29.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.84$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+10.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.50$}
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.37$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-4.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+14.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.92$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.14$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-3.77$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.85$}
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.75$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.66$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.68$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+16.14$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.26$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.60$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.24$}
@@
$\Delta$(PatMD, Vanilla-MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.56$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.56$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.60$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.62$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.67$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.65$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+22.30$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+30.81$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.39$}

### v3

@@ 3.2.1. Bipartite Candidate Retrieval
Specifically, we use the target sample’s image vector $v_{\text{test}}^{\text{img}}$ and text vector $v_{\text{test}}^{\text{txt}}$ to separately query the subsets of samples corresponding to benign and harmful ground-truth labels. Within each labeled subset, we retrieve the top-$N$ most similar candidates from the image modality and the top-$N$ from the text modality. Subsequently, we take the union of all these retrieved candidates to form an initial candidate pool. This pool, containing samples that offer a balanced perspective on potential risks, is then passed to the reranking part for a more refined similarity assessment and reranking.
@@ 3.2.2. Hierarchical Harm Tree-based Reranking
The product term $\prod_{i=1}^{k}M_{i}$ ensures that the bonus $\gamma$ is awarded only if all levels from $L_{1}$ to $L_{k}$ have successfully matched, thus enforcing the top-down sequential mechanism. Consequently, the tree similarity score for a candidate, denoted as $S_{\mathcal{T}}$, is the hierarchical bonuses accumulated from all continuously matched levels (top-to-down).
Figure 3 illustrates this advantage. Consider two hateful memes sharing the same derogatory logic (i.e., exploiting the slang “vegetable”). Due to distinct surface features, conventional content matching fails to link them, yielding a low similarity score of 0.55. Conversely, PatMD hierarchically deconstructs them to capture implicit rhetorical analogies. By accumulating structural similarities across $L_{1}$$L_{4}$ alongside hierarchical bonuses (+0.4 weight score), PatMD bridges the semantic gap, boosting the final similarity score to 0.95 for accurate retrieval.
@@
where $\alpha$ is a hyperparameter that balances the importance of the image and text modalities, and $S_{\mathcal{T}}$ represents the score derived from the structured harm matching.
All candidates are sorted in descending order based on their $S_{\text{final}}$ score, and the top-$K$ samples are selected. The corresponding misjudgment risk patterns of these selected samples are subsequently used to guide the follow-up detection.
@@ 3.3. Pattern-Augmented Reasoning for Detection
We present pattern-augmented prompt construction, where the prompt is built upon a structured template designed to orchestrate the MLLM’s analytical workflow. The template first grounds the MLLM by clearly stating the classification task and providing an objective definition of the target harm category adopted from (Lin et al., 2025). Its core component is the Critical Reflection References section, which contains the retrieved misjudgment patterns categorized based on their ground-truth labels.
Patterns associated with harmful memes are aggregated and framed as Miss Reasons to alert the MLLM to subtle harmful patterns it might otherwise overlook. Conversely, patterns from benign memes are presented as False Alarm Reasons to prevent the MLLM from misclassifying non-hateful content. To ensure this guidance is used for critical reflection rather than simple imitation, the template also enforces a structured reasoning procedure that compels the model to form an independent analysis before consulting the provided support. Through this strategy, a unique and highly relevant prompt is dynamically generated for each target sample (See Appendix C Figure 10 for detailed prompts).
@@ 4. Experiment Design
  • RQ2: What is the contribution of components to the harmful meme detection? Since PatMD’s framework follows the stages of the RAG paradigm, we aim to evaluate how PatMD outperforms the original RAG’s strategy, by comparing the performance of PatMD and variants of removing components.
  • RQ3: What is the performance of PatMD on unseen and adversarial memes? We aim to evaluate the generality and robustness of our approach PatMD under out-of-distribution and perturbed scenarios.
@@ 4.1. Datasets
PromptHarm (Cao et al., 2022) 64.05 51.57 69.30 67.81 61.88 57.89 58.51 55.08 61.71 53.44 63.28 57.16
Pro-Cap (Cao et al., 2023) 52.40 51.80 61.56 58.57 48.56 47.15 54.90 54.62 57.06 56.83 54.90 53.79
CapAlign (Ji et al., 2024) 63.70 51.60 68.60 66.81 61.51 54.76 58.08 52.67 60.77 51.04 62.53 55.38
MemeAgent (Lin et al., 2025) 68.40 62.73 73.10 73.04 62.18 58.96 70.11 69.97 63.23 61.25 67.40 65.19
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.94$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+17.95$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.53$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.20$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.10$}
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.40$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-2.61$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.00$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.17$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.49$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.71$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+12.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.48$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.72$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$}
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.82$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+15.78$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.18$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-1.88$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+21.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+29.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+8.84$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.24$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+10.99$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.50$}
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.37$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-4.40$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+11.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+14.59$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.11$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.92$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+20.70$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.14$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-3.77$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+9.91$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.85$}
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.35$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.90$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+3.75$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.66$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.19$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+13.68$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+16.14$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.26$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.60$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.89$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+5.24$}
@@
$\Delta$(PatMD, MLLM) $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+2.56$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+0.56$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.60$} $\color[rgb]{0,0.19921875,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0}-0.62$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.67$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+4.65$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+22.30$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+30.81$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.52$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+1.54$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+6.49$} $\color[rgb]{0.6015625,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.6015625,0,0.3984375}+7.39$}